OpenStudy (apoorvk):

Three friends Roger, Djoko and Rafael (:p) went to watch a tournament of tennis on a certain day of the event, when a total of 7 matches were being held (no match-timings coinciding with any other). All the 3 friends can either watch a match or loiter around in the arena with the probability ratio (3:1) On entering the arena, all 3 went their separate ways. Find the probability of none of them having witnessed any particular match together.

OpenStudy (anonymous):

:( I can't solve it, @foolformath HELP.

OpenStudy (apoorvk):

nahii.. ISHAAN, ye naaahhiii ho saktaa!! :p :D

OpenStudy (anonymous):

I can't also.

OpenStudy (anonymous):

lol @mr.math @dumbcow

OpenStudy (apoorvk):

@myininaya

OpenStudy (apoorvk):

help myininaya, please!!!

OpenStudy (anonymous):

should be 1 in 448

OpenStudy (anonymous):

How?

OpenStudy (apoorvk):

now explanation will be appreciated. :)

OpenStudy (apoorvk):

am getting stuck somewhere. and i can't untangle my mind.. :/

OpenStudy (anonymous):

Change my answer. 27 in 448 chance that they will all meet, 136 in 448 chance that a and b will meet, 136 in 448 chance that b and c will meet, 136 in 448 chance that a and c will meet. 3*136- 2*27= 354 in 448 chance that they will meet at all. Therefore 94 in 448 (you can simplify it) chance that they will not meet at all.

OpenStudy (anonymous):

Timings of the match doesn't coincide with eachother, so when they enter the stadium they must have only one match to watch, lets they don't watch it together. So, there must be three cases either one of them watches the match or two or none. \[\text{_________________________________________________________________________________}\]Now help me do the math.

OpenStudy (lgbasallote):

no one's calling for lgbaaaa :( well i dunno this either...

OpenStudy (mani_jha):

For each of the three persons, there are two choices - watch the match or not, for each of the seven matches. So, total no of ways in which each one completes the day is 2X2X2..=2^7. Is this right so far?

OpenStudy (mani_jha):

The total no of choices for the 3 people would be 3X2^7. Atleast I found out the denominator of my fraction..:P

OpenStudy (apoorvk):

right so far mani, and i seem to be seeing it clearer now.

OpenStudy (apoorvk):

unhnn....the choices for the three would get multiplied rendering your denominator 2^21, since it is an 'OR' case, where each member may watch a match or not, and each person's choice being an independent event.

OpenStudy (apoorvk):

'OR'

OpenStudy (mani_jha):

oh yeah. you're right. Now let me think how to use the given probability ratio.

OpenStudy (mani_jha):

@ch2831, where did you get the number 448 from?

OpenStudy (anonymous):

The names have been badly chosen :P

OpenStudy (anonymous):

4*4*4*7. Isnt it 448?

OpenStudy (mani_jha):

Yeah, it is. Please explain why you chose that number. To be honest, I didn't understand your answer.

OpenStudy (dumbcow):

hmm i'm getting a really small probability here --> (5/32)^7 each match is an independent event so you would multiply the probabilities hence raising to 7th power each match has 4 possiblilties None watch: --> (1/4)^3 = 1/64 1 watches: --> 3*(3/4)*(1/4)^2 = 9/64 2 watches --> 3*(3/4)^2*(1/4) = 27/64 All watch --> (3/4)^3 = 27/64 so probability less than 2 watch any given match is 10/64 = 5/32

OpenStudy (anonymous):

Oh so this is how the probability ratio is used

OpenStudy (mani_jha):

Oh...This topic really isn't made for me :=(

OpenStudy (dumbcow):

i believe that is right, the probability you watch is 3 times greater than not watching

OpenStudy (apoorvk):

@dumbcow .. yayy!!!!!!!!!! you did it i guess!!! yeah he took the ratio thing right too, ishaan!

OpenStudy (apoorvk):

@arcticf0x , the names have been *purposefully* chosen. :p

OpenStudy (anonymous):

Nice solution :-) #I need to study probability :/